OpenStudy (anonymous):

solve for the variable. -x/9 = 5/3

jimthompson5910 (jim_thompson5910):

Hint cross multiply to get -3x = 45

OpenStudy (anonymous):

so that would be -15 but before that you do the reciprocal of 5/3 right?

jimthompson5910 (jim_thompson5910):

correct -3x = 45 x = 45/(-3) x = -15

jimthompson5910 (jim_thompson5910):

not sure what you mean by the part "but before that you do the reciprocal of 5/3 right"

OpenStudy (anonymous):

you switched the 5& 3. so the three could be on top and the five in the bottom\[\frac{ 5 }{ 3} = \frac{ 3 }{ 5 }\]

jimthompson5910 (jim_thompson5910):

oh I started with 

jimthompson5910 (jim_thompson5910):

then I cross multiplied by multiplying like this